# Efficient tomography of microwave photonic cluster states

### v2

@@ E.1 Mode function of the photonic qubit
First, the mode function of the photonic qubit needs to be measured because we obtain the quadrature values by amplifying the quadrature using a JPA and calculating the overlap integral between the amplified waveform and the mode function. We use the pulse sequence shown in Fig. 16(a) to generate a photonic qubit in the superposition state $\lvert+\rangle\coloneqq(\lvert 0\rangle+\lvert 1\rangle)/\sqrt{2}$ and measure its averaged waveform. To obtain a complex-valued waveform, we perform another measurement with the phase of the $\pi_{ge}/2$ pulse shifted by $\pi/2$. This generates a photonic qubit in $\lvert+i\rangle\coloneqq(\lvert 0\rangle+i\lvert 1\rangle)/\sqrt{2}$, whose averaged waveform is phase-shifted by $\pi/2$. We also measure the waveforms of photonic qubits in $\lvert-\rangle\coloneqq(\lvert 0\rangle-\lvert 1\rangle)/\sqrt{2}$ and $\lvert-i\rangle\coloneqq(\lvert 0\rangle-i\lvert 1\rangle)/\sqrt{2}$ and subtract them from those of |+delimited-|$\lvert$+$\rangle$| + ⟩ and $\lvert+i\rangle$, respectively, to remove any offset and background signal in the waveforms.
Figure 16(b) shows the demodulated complex waveform obtained with and without the amplification by the JPA. To phase-insensitively amplify the signal using the phase-sensitive mode of the JPA, we add the results of two sets of measurements with the phase of the JPA pump set to 0 and $\pi$. This results in a phase-insensitive amplitude gain of $\sqrt{S}+1/\sqrt{S}$, where $S$ is the squeezing factor of the JPA. A squeezing factor of $15.710\log_{10}S=$ 15.7 dB is obtained from the measured waveforms. The amplification by the JPA delays but does not significantly deform the waveform, which suggests that the gain bandwidth of the JPA is sufficient for the photonic qubit. The waveform measured with the amplification by the JPA is used as the mode function when calculating the quadrature values in the tomography experiments.
@@ E.2 Pauli tomography using quadrature measurements
[…]
@@
where $\hat{\Pi}\coloneqq\lvert 0\rangle\!\langle 0\rvert+\lvert 1\rangle\!\langle 1\rvert$ is the projection operator. By taking the expectation values of these equations, the conversion formula shown in Eqs. (7a)=̄(7c) is obtained. The expectation values of the Pauli observables can therefore be estimated as
@@
[…]
@@
[…]
@@ E.3 Correcting for the measurement inefficiency
[…]
@@
The measurement efficiency $\eta$ is obtained as the ratio between $\mkern 1.5mu\overline{\mkern-1.5mun\mkern-1.5mu}\mkern 1.5mu_{\mathrm{coh}}$ and the actual mean photon number of the coherent pulse, which is calculated using the attenuation of the input line estimated by the post-selected resonator spectroscopy described in Ref. 42. The amplitude of the coherent pulse is varied to confirm that the pulse is not saturating the JPA. Figure 17(b) shows the result and the linear fit, which gives a measurement efficiency of $\eta=0.391\pm 0.004$.
@@ E.4 Pauli tomography of multiple photonic qubits
Here, we introduce the multi-qubit generalization of the conversion formula between the quadrature and Pauli measurements. To derive the formula, let us split Eqs. (38a)=̄(38c) into two steps. The first step converts the quadrature moments to the “$Z$-shifted Pauli basis” as
@@
is the conversion matrix. Using the notation defined in Eqs. (6a)=̄(6c) for the quadrature moments and the $Z$-shifted Pauli basis defined as
@@
We compare the reconstructed density matrix with a numerical model with uniform photon loss and dephasing probabilities across all qubits. The photon loss error of the cluster state can be evaluated by measuring the mean excitation $(1-\langle\hat{Z}_{s}\rangle)/2$ of each photonic qubit, which is affected by the photon loss but not by the dephasing error. Figure 18(a) shows the measured mean excitations and the fit, which gives an average photon loss probability of 9.8%. The dephasing error can be evaluated by measuring the expectation values of the stabilizer operators of the cluster state, which are affected by both the photon loss and dephasing errors. For this, we first need to align the phase origins of the photonic qubits with the phase origins of the quadrature measurements. This is done by numerically rotating the $s$th photonic qubit by an angle of $-\!\arctan(\langle\hat{Z}_{s-1}\hat{Y}_{s}\hat{Z}_{s+1}\rangle/\langle\hat{Z}_{s-1}\hat{X}_{s}\hat{Z}_{s+1}\rangle)$, which maximizes the stabilizer $\langle\hat{Z}_{s-1}\hat{X}_{s}\hat{Z}_{s+1}\rangle$. This correction has been applied to all data presented in this work. Then, we obtain the stabilizers shown in Fig. 18(b), which give an average dephasing error equivalent to a phase flip probability of 4.6%. Note that the photon loss error has a weaker effect on the first and last stabilizers because they contain fewer $\hat{Z}$ operators than the other stabilizers.
Figures 19(a)=̄(c) show the density matrices of an ideal five-qubit cluster state, the numerical model with uniform error probabilities, and the generated cluster state, respectively. Whereas the complex arguments of the three density matrices match closely, the absolute values of the matrix elements of the experimentally generated cluster state deviate significantly from the ideal value of $0.031252^{-5}=$ 0.03125 because they are affected by the photon loss and dephasing errors. The fidelity of the cluster state to the numerical model is $\mathcal{F}=0.970$, which suggests that the numerical model captures the dominant error processes. The fidelity of the cluster state to the ideal state is $\mathcal{F}=0.616\pm 0.006$, which is significantly larger than the threshold of 0.5 for genuine multipartite entanglement.
@@
Here, we present a simplified reformulation of the MPO reconstruction procedure introduced in Ref. 34. This reformulation provides an explicit formula for the reconstructed MPO in terms of the measured local correlations, which is used in Sec. IV.2 to construct an initial guess for the least-squares fitting. It is also used to show that five-qubit local correlations are required to reconstruct a linear cluster state, despite the fact that three-qubit local correlations are sufficient for reconstructing the vast majority of MPOs with the same bond dimension.
@@ G.1 Using up to three-qubit local correlations
Figure 20 shows the tensor-network representation of the reconstructed MPO given by Eqs. (54a)=̄(54d). To verify this solution, let us express the measured two- and three-qubit local correlations of the true state as
@@
Since a randomly chosen matrix almost always has full rank, an MPO with a bond dimension $D\leq 4$ almost always satisfies the reconstructibility condition. Therefore, one might expect that a linear cluster state, which has a bond dimension of $D=4$, can be reconstructed from three-qubit local correlations. However, it turns out that linear cluster states do not satisfy the reconstructibility condition. This can be seen by showing that Eq. (54c) has no solution as follows. Because the stabilizer operators of a linear cluster state are given by Eqs. (11a)=̄(11c), the only non-zero two- and three-qubit local correlations of a linear cluster state are
@@ H.2 Estimating the bond dimension
It follows from the above procedure that the bond dimension between the $(s+1)$th and $(s+2)$th qubits can be experimentally determined by measuring the four-qubit correlation matrix $B_{s}$ and counting the number of positive singular values. However, because the measured four-qubit correlation matrix contains statistical noise, its singular values need to be significantly larger than their statistical uncertainties to qualify as being positive. The statistical uncertainty of a measured four-qubit correlation matrix $\mkern 1.5mu\overline{\mkern-1.5muB\mkern-1.5mu}\mkern 1.5mu$ can be propagated to the singular values by calculating the partial derivative of a singular value $\sigma_{n}$ by the $(i,j)$ matrix element as
$\frac{\partial\sigma_{n}}{\partial\mkern 1.5mu\overline{\mkern-1.5muB\mkern-1.5mu}\mkern 1.5mu_{i,j}}=U_{i,n}V^{*}_{j,n},$ (81)
@@ I.2 Transformation into the standard form
Figure 24: Tensor-network representation of the unnormalized post-measurement state $\hat{\rho}^{\prime\{m_{s}\}}$, where ${\{m_{s}\}}$ denotes the set of measurement outcomes $m_{s}\in\{+1,-1\}$ obtained by performing a local projective measurement $\hat{\Pi}_{s}^{\pm 1}$ on every qubit of an MPO except the $r$th and $r^{\prime}$th qubits.
A projective measurement of a qubit is defined by a pair of orthogonal projection operators $\hat{\Pi}^{\pm 1}$, where ±$\pm 1$ are the outcomes of the measurement. If this measurement is performed on the $s$th qubit in an $N$-qubit state $\hat{\rho}$, the probability of obtaining each outcome is
@@
Now suppose that a local projective measurement $\hat{\Pi}_{s}^{\pm 1}$ is performed on every qubit of an MPO $\hat{\mathsf{A}}_{s}$ except the $r$th and $r^{\prime}$th qubits. The unnormalized post-measurement state $\hat{\rho}^{\prime\{m_{s}\}}$, where ${\{m_{s}\}}$ denotes the set of measurement outcomes $m_{s}\in\{+1,-1\}$, can be calculated by evaluating the tensor network shown in Fig. 24. Similarly to Fig. 7(b), the partial derivative by a parameter of the MPO $(A^{(i_{s})}_{s})_{x,y}$ can be calculated by removing the tensor $\hat{\mathsf{A}}_{s}$ from the diagram. Using the unnormalized post-measurement state $\hat{\rho}^{\prime\{m_{s}\}}$, the outcome probability can be expressed as
$P(\{m_{s}\})=\operatorname{Tr}\left[\hat{\rho}^{\prime\{m_{s}\}}\right]$ (100)
@@
$\hat{\rho}^{\{m_{s}\}}=\frac{\hat{\rho}^{\prime\{m_{s}\}}}{\operatorname{Tr}\left[\hat{\rho}^{\prime\{m_{s}\}}\right]}.$ (101)
@@
[…]
@@
$\mathcal{N}_{\mathrm{le}}(\hat{\rho})\coloneqq\sum_{\{m_{s}\}}P(\{m_{s}\})\,\mathcal{N}(\hat{\rho}^{\{m_{s}\}}),$ (105)
given a measurement basis $\hat{\Pi}_{s}^{\pm 1}$ for every qubit except the two unmeasured ones at $s=r$ and $r^{\prime}$. Note that the original definition of the localizable entanglement is the maximum of $\mathcal{N}_{\mathrm{le}}(\hat{\rho})$ over all possible measurement bases $\{\hat{\Pi}_{s}^{\pm 1}\}_{s\neq r,r^{\prime}}$, which is computationally costly to determine [24]. Because we do not perform the maximization, the definition used in this work is only a lower bound of the original definition. Nevertheless, the optimal set of measurement bases for the ideal linear cluster state can be used to obtain a nearly maximal $\mathcal{N}_{\mathrm{le}}(\hat{\rho})$ for a state which is close to an ideal cluster state. To calculate the localizable entanglement of the generated cluster states, we use the $\hat{X}_{s}$ measurement basis for all qubits between the two unmeasured qubits and the $\hat{Z}_{s}$ basis for all the other qubits.
To calculate the uncertainty of the localizable entanglement, Eq. (105) can be rewritten using the unnormalized post-measurement states $\hat{\rho}^{\prime\{m_{s}\}}$ as
$\mathcal{N}_{\mathrm{le}}(\hat{\rho})=\sum_{\{m_{s}\}}\mathcal{N}(\hat{\rho}^{\prime\{m_{s}\}}).$ (106)
@@
$\mathcal{C}_{\mathrm{le}}(\hat{\rho})=\sum_{\{m_{s}\}}\mathcal{C}(\hat{\rho}^{\prime\{m_{s}\}}),$ (107)
where $\mathcal{C}(\hat{\rho})$ denotes the concurrence of a two-qubit state $\hat{\rho}$.

### v3

@@ E.1 Mode function of the photonic qubit
First, the mode function of the photonic qubit needs to be measured because we obtain the quadrature values by amplifying the quadrature using a JPA and calculating the overlap integral between the amplified waveform and the mode function. We use the pulse sequence shown in Fig. 16(a) to generate a photonic qubit in the superposition state $\lvert+\rangle\coloneqq(\lvert 0\rangle+\lvert 1\rangle)/\sqrt{2}$ and measure its averaged waveform. To obtain a complex-valued waveform, we perform another measurement with the phase of the $\pi_{ge}/2$ pulse shifted by $\pi/2$. This generates a photonic qubit in $\lvert+i\rangle\coloneqq(\lvert 0\rangle+i\lvert 1\rangle)/\sqrt{2}$, whose averaged waveform is phase-shifted by $\pi/2$. We also measure the waveforms of photonic qubits in $\lvert-\rangle\coloneqq(\lvert 0\rangle-\lvert 1\rangle)/\sqrt{2}$ and $\lvert-i\rangle\coloneqq(\lvert 0\rangle-i\lvert 1\rangle)/\sqrt{2}$ and subtract them from those of $\lvert$$\rangle$ and $\lvert+i\rangle$, respectively, to remove any offset and background signal in the waveforms. Note that the complex-valued waveform can alternatively be obtained, without phase-shifting the photon, by applying a Hilbert transform to the measured real-valued waveform, which is at the intermediate frequency (IF) of the down-converting mixer.
Figure 16(b) shows the demodulated complex waveform obtained with and without the amplification by the JPA. To phase-insensitively amplify the signal using the phase-sensitive mode of the JPA, we add the results of two sets of measurements with the phase of the JPA pump set to 0 and $\pi$. This results in a phase-insensitive amplitude gain of $\sqrt{S}+1/\sqrt{S}$, where $S$ is the squeezing factor of the JPA. A squeezing factor of $10\log_{10}S=15.7$ dB is obtained from the measured waveforms. The amplification by the JPA delays but does not significantly deform the waveform, which suggests that the gain bandwidth of the JPA is sufficient for the photonic qubit. The waveform measured with the amplification by the JPA is used as the mode function when calculating the quadrature values in the tomography experiments.
@@ E.2 Pauli tomography using quadrature measurements
To derive the conversion formula between the quadrature measurements and the Pauli measurements, let us project the quadrature operators onto the $\{\lvert 0\rangle,\lvert 1\rangle\}$ subspace as
@@
where $\hat{\Pi}\coloneqq\lvert 0\rangle\!\langle 0\rvert+\lvert 1\rangle\!\langle 1\rvert$ is the projection operator. By taking the expectation values of these equations, the conversion formula shown in Eqs. (7a)(7c) is obtained. The expectation values of the Pauli observables can therefore be estimated as
@@
$\mkern 1.5mu\overline{\mkern-1.5muZ_{s}\mkern-1.5mu}\mkern 1.5mu$ $=2-\mkern 1.5mu\overline{\mkern-1.5muq_{s}^{2}\mkern-1.5mu}\mkern 1.5mu-\mkern 1.5mu\overline{\mkern-1.5mup_{s}^{2}\mkern-1.5mu}\mkern 1.5mu,$ (40c)
@@
$\operatorname{SE}\left[\mkern 1.5mu\overline{\mkern-1.5muZ_{s}\mkern-1.5mu}\mkern 1.5mu\right]$ $=\sqrt{\operatorname{Var}\left[\mkern 1.5mu\overline{\mkern-1.5muZ_{s}\mkern-1.5mu}\mkern 1.5mu\right]}$ $=\sqrt{\operatorname{Var}\bigl[\mkern 1.5mu\overline{\mkern-1.5muq_{s}^{2}\mkern-1.5mu}\mkern 1.5mu\bigr]+\operatorname{Var}\bigl[\mkern 1.5mu\overline{\mkern-1.5mup_{s}^{2}\mkern-1.5mu}\mkern 1.5mu\bigr]},$ (41c)
@@ E.3 Correcting for the measurement inefficiency
$\begin{bmatrix}1\\ \mkern 1.5mu\overline{\mkern-1.5muX_{s}\mkern-1.5mu}\mkern 1.5mu\\ \mkern 1.5mu\overline{\mkern-1.5muY_{s}\mkern-1.5mu}\mkern 1.5mu\\ \mkern 1.5mu\overline{\mkern-1.5muZ_{s}\mkern-1.5mu}\mkern 1.5mu\end{bmatrix}=\mathcal{E}_{\mathrm{ad}}^{-1}\begin{bmatrix}1\\ \mkern 1.5mu\overline{\mkern-1.5muX_{s}^{\prime}\mkern-1.5mu}\mkern 1.5mu\\ \mkern 1.5mu\overline{\mkern-1.5muY_{s}^{\prime}\mkern-1.5mu}\mkern 1.5mu\\ \mkern 1.5mu\overline{\mkern-1.5muZ_{s}^{\prime}\mkern-1.5mu}\mkern 1.5mu\end{bmatrix}=\begin{bmatrix}1\\ \eta^{-1/2}\mkern 1.5mu\overline{\mkern-1.5muX_{s}^{\prime}\mkern-1.5mu}\mkern 1.5mu\\ \eta^{-1/2}\mkern 1.5mu\overline{\mkern-1.5muY_{s}^{\prime}\mkern-1.5mu}\mkern 1.5mu\\ 1-\eta^{-1}(1-\mkern 1.5mu\overline{\mkern-1.5muZ_{s}^{\prime}\mkern-1.5mu}\mkern 1.5mu)\end{bmatrix}$, (42)
@@
The measurement efficiency $\eta$ is obtained as the ratio between $\mkern 1.5mu\overline{\mkern-1.5mun\mkern-1.5mu}\mkern 1.5mu_{\mathrm{coh}}$ and the actual mean photon number of the coherent pulse. The amplitude of the coherent pulse is varied to confirm that the pulse is not saturating the JPA. Figure 17(b) shows the result and the linear fit, which gives a measurement efficiency of $\eta=0.391\pm 0.004$. Here, the actual mean photon number of the coherent pulse is calculated using the attenuation of the input line, which is estimated by measuring the qubit dephasing while applying a weak coherent drive at the same frequency as the pulse [30].
Note that our estimation of the measurement efficiency places the reference plane at the output port of the sample. This means that the photonic states reconstructed in this work reflect the imperfections of the photon emission process, such as the energy relaxation of the $\lvert f\rangle$ state during the photon emission and the infidelity of the $\pi_{ef}$ pulse.
@@ E.4 Pauli tomography of multiple photonic qubits
Here, we introduce the multi-qubit generalization of the conversion formula between the quadrature and Pauli measurements. To derive the formula, let us split Eqs. (39a)(39c) into two steps. The first step converts the quadrature moments to the “$Z$-shifted Pauli basis” as
@@
is the conversion matrix. Using the notation defined in Eqs. (6a)(6c) for the quadrature moments and the $Z$-shifted Pauli basis defined as
@@
We compare the reconstructed density matrix with a numerical model with uniform photon loss and dephasing probabilities across all qubits. The photon loss error of the cluster state can be evaluated by measuring the mean excitation $(1-\langle\hat{Z}_{s}\rangle)/2$ of each photonic qubit, which is affected by the photon loss but not by the dephasing error. Figure 18(a) shows the measured mean excitations and the fit, which gives an average photon loss probability of 9.8%. The dephasing error can be evaluated by measuring the expectation values of the stabilizer operators of the cluster state, which are affected by both the photon loss and dephasing errors. For this, we first need to align the phase origins of the photonic qubits with the phase origins of the quadrature measurements. This is done by numerically rotating the $s$th photonic qubit by an angle of $\arctan(\langle\hat{Z}_{s-1}\hat{Y}_{s}\hat{Z}_{s+1}\rangle/\langle\hat{Z}_{s-1}\hat{X}_{s}\hat{Z}_{s+1}\rangle)$, which maximizes the stabilizer $\langle\hat{Z}_{s-1}\hat{X}_{s}\hat{Z}_{s+1}\rangle$. This correction has been applied to all data presented in this work. Then, we obtain the stabilizers shown in Fig. 18(b), which give an average dephasing error equivalent to a phase flip probability of 4.6%. Note that the photon loss error has a weaker effect on the first and last stabilizers because they contain fewer $\hat{Z}$ operators than the other stabilizers.
Figures 19(a)(c) show the density matrices of an ideal five-qubit cluster state, the numerical model with uniform error probabilities, and the generated cluster state, respectively. Whereas the complex arguments of the three density matrices match closely, the absolute values of the matrix elements of the experimentally generated cluster state deviate significantly from the ideal value of $2^{-5}=0.03125$ because they are affected by the photon loss and dephasing errors. The fidelity of the cluster state to the numerical model is $\mathcal{F}=0.970$, which suggests that the numerical model captures the dominant error processes. The fidelity of the cluster state to the ideal state is $\mathcal{F}=0.616\pm 0.006$, which is significantly larger than the threshold of 0.5 for genuine multipartite entanglement.
@@
Here, we present a simplified reformulation of the MPO reconstruction procedure introduced in Ref. 8. This reformulation provides an explicit formula for the reconstructed MPO in terms of the measured local correlations, which is used in Sec. IV.2 to construct an initial guess for the least-squares fitting. It is also used to show that five-qubit local correlations are required to reconstruct a linear cluster state, despite the fact that three-qubit local correlations are sufficient for reconstructing the vast majority of MPOs with the same bond dimension.
@@ G.1 Using up to three-qubit local correlations
Figure 20 shows the tensor-network representation of the reconstructed MPO given by Eqs. (55a)(55d). To verify this solution, let us express the measured two- and three-qubit local correlations of the true state as
@@
Since a randomly chosen matrix almost always has full rank, an MPO with a bond dimension $D\leq 4$ almost always satisfies the reconstructibility condition. Therefore, one might expect that a linear cluster state, which has a bond dimension of $D=4$, can be reconstructed from three-qubit local correlations. However, it turns out that linear cluster states do not satisfy the reconstructibility condition. This can be seen by showing that Eq. (55c) has no solution as follows. Because the stabilizer operators of a linear cluster state are given by Eqs. (11a)(11c), the only nonzero two- and three-qubit local correlations of a linear cluster state are
@@ H.2 Estimating the bond dimension
It follows from the above procedure that the bond dimension between the $(s+1)$th and $(s+2)$th qubits can be experimentally determined by measuring the four-qubit correlation matrix $B_{s}$ and counting the number of positive singular values. However, because the measured four-qubit correlation matrix contains statistical noise, its singular values need to be significantly larger than their statistical uncertainties to qualify as being positive. The statistical uncertainty of a measured four-qubit correlation matrix $\mkern 1.5mu\overline{\mkern-1.5muB\mkern-1.5mu}\mkern 1.5mu$ can be propagated to the singular values to the first order by calculating the partial derivative of a singular value $\sigma_{n}$ by the $(i,j)$ matrix element as
$\frac{\partial\sigma_{n}}{\partial\mkern 1.5mu\overline{\mkern-1.5muB\mkern-1.5mu}\mkern 1.5mu_{i,j}}=U^{*}_{i,n}V_{j,n},$ (82)
@@ I.2 Transformation into the standard form
[…]
A projective measurement of a qubit is defined by a pair of orthogonal projection operators $\hat{\Pi}^{\pm 1}$, where $\pm 1$ are the outcomes of the measurement. If this measurement is performed on the $s$th qubit in an $N$-qubit state $\hat{\rho}$, the probability of obtaining each outcome is
@@
[…]
@@
[…]
@@
where $\rho{\ulcorner}$ denotes the partial transpose of $\hat{\rho}$ in the $\{\lvert 00\rangle,\lvert 01\rangle,\lvert 10\rangle,\lvert 11\rangle\}$ basis,
@@
[…]
@@
[…]